# TNPAR: Topological Neural Poisson Auto-Regressive Model for Learning Granger Causal Structure from Event Sequences

### v1

@@ Abstract
Learning Granger causality from event sequences is a challenging but essential task across various applications. Most existing methods rely on the assumption that event sequences are independent and identically distributed (i.i.d.). However, this i.i.d. assumption is often violated due to the inherent dependencies among the event sequences. Fortunately, in practice, we find these dependencies can be modeled by a topological network, suggesting a potential solution to the non-i.i.d. problem by introducing the prior topological network into Granger causal discovery. This observation prompts us to tackle two ensuing challenges: 1) how to model the event sequences while incorporating both the prior topological network and the latent Granger causal structure, and 2) how to learn the Granger causal structure. To this end, we devise a two-stage unified topological neural Poisson auto-regressive model. During the generation stage, we employ a variant of the neural Poisson process to model the event sequences, considering influences from both the topological network and the Granger causal structure. In the inference stage, we formulate an amortized inference algorithm to infer the latent Granger causal structure. We encapsulate these two stages within a unified likelihood function, providing an end-to-end framework for this task.
1 Introduction
@@
Causal discovery from multi-type event sequences plays a pivotal role in a multitude of applications. In the realm of intelligent operation and maintenance, for instance, identifying the causal structure behind alarm data can expedite the pinpointing of critical fault events and the location of root causes Vuković and Thalmann (2022). In social analysis, understanding the causal structure behind users’ behavior sequences can inform the development of effective advertising and marketing strategies Chen et al. (2020); Wu et al. (2022).
@@
A range of methods has been proposed for the causal discovery of multi-type event sequences. They fall into two main categories: 1) constraint-based methods, which utilize independence-based tests or measures to ascertain the causal structure between event types, and 2) point process-based methods, which employ point processes to model the generation process of event sequences. While the former, including methods such as PCMCI (PC with Momentary Conditional Independence test) Runge (2020) and transfer entropy-based method Chen et al. (2020); Mijatovic et al. (2021), is founded on strict assumptions of the causal mechanism assumptions, the latter, like Hawkes process-based methods Xu et al. (2016); Zhou et al. (2013); Eichler et al. (2017) and neural point process models Shchur et al. (2021); Chen et al. (2021), is grounded in the concept of Granger causality Granger (1969). Given the weaker assumption of Granger causality, identifying the Granger causal structure just by assessing if a series is predictive of another, point process-based methods often find preference in real-world applications Shojaie and Fox (2022) and are therefore the focal point of this work.
@@
Figure 1: An illustrative example of the topological event sequences generated by a mobile network in an operation and maintenance scenario. In this context, $\{n_{1},n_{2},n_{3}\}$ represent the network elements, and $\{v_{1},v_{2},v_{3}\}$ denote the event types of the alarms. The term $\mathrm{P}_{v_{1}}(t_{1})$ stands for the distribution of $v_{1}$ alarm at $t_{1}$ timestamp, and so on. Black arrows highlight the correct causal edges, whereas the red arrow represents an incorrect edge. (a) Generation process of topological event sequences. The event distribution varies with time and is influenced by both a topological network (connected by the solid lines) and a Granger causal structure (depicted in part (b)). (b) Ground truth of the Granger causal structure. (c) Granger causal structure learned under i.i.d. assumption.
@@
However, these existing point process-based methods are significantly dependent on the i.i.d assumption (that the event sequences are independent and identically distributed) to model the generation process of event sequences. In reality, this assumption often falls short due to the complex dependencies that permeate sequences. Consequently, the performance of these methods often leaves much to be desired in real-world applications. For instance, non-i.i.d. alarm event sequences in an operation and maintenance scenario, as illustrated in Fig. 1, highlight this issue. These event sequences are produced by network elements interconnected by a topological network, indicating the event sequences of different elements are dependent on each other, thereby violating the i.i.d. assumption. In this illustrated scenario, the event type $v_{2}$ in $n_{2}$ is only caused by the event types in topologically connected elements (specifically, $v_{1}$ in $n_{2}$). Existing methods would incorrectly identify the edge $v_{3}\rightarrow v_{2}$ because they erroneously treat event sequences on different nodes as independent. By incorporating the topological network, we can correctly learn that $v_{2}$ is actually caused by $v_{1}$ and discern the true causal structure. Therefore, it is vital to exploit the topological network to model the dependencies among non-i.i.d. event sequences and accurately recover the causal structure from them.
@@
This gives rise to two ensuing challenges: 1) how to explicitly model the generation process of topological event sequences, incorporating both the prior topological network and the latent causal structure, and 2) how to learn the latent Granger causal structure from topological event sequences. To tackle these challenges, Cai et al. Cai et al. (2022) propose the topological Hawkes process (THP), which extends the time-domain convolution property of the Hawkes process to the time-graph domain. Nevertheless, THP’s applicability in real-world scenarios remains limited due to the complex generation process and distribution of real events, which are often too intricate to be accurately represented by the Hawkes process. Consequently, Hawkes process-based methods often suffer from oversimplified generation functions and specified distributions. Hence, it is crucial to devise a distribution-free generation model that can incorporate both the topological event sequences and the latent Granger causal structure.
@@
In response to these challenges, we develop a unified Topological Neural Poisson Auto-Regressive (TNPAR) model, comprising a generation process and an inference process. In the generation process, we introduce a variant of the Poisson Auto-Regressive model to depict the generation process of topological event sequences. By extending the Auto-Regressive model with Poisson Process and incorporating both the topological network and the Granger causal structure, TNPAR effectively overcomes the non-i.i.d. problem. By representing the distribution using a series of Poisson processes with varying parameters, TNPAR displays enhanced flexibility compared to existing methods that employ a single fixed distribution for the entire sequence. For the inference process, inspired by the generative model that treats the causal structure as a latent variable Li et al. (2021), we design an amortized inference algorithm Zhang et al. (2018); Gershman and Goodman (2014) to learn the Granger causal structure. By integrating these generation and inference procedures, we establish a unified likelihood function for our TNPAR model, which can be optimized within an end-to-end paradigm.
@@
    Our method incorporates the prior topological network into the generation process, offering a solution to the non-i.i.d. challenge in causal discovery.
@@
    By treating the causal structure as a latent variable, we devise an amortized inference method to identify the causal structure among the event types.
2 Related work
This section presents the related work on Granger causal discovery from event sequences. We have also provided a comparison of our work with related methods in Appendix A.
2.1 Temporal Point Process
@@
Temporal point processes are stochastic processes utilized to model event sequences. Current work based on temporal point processes can be bifurcated into two main categories: statistical point processes and neural point processes. On the one hand, statistical point processes rely on parametric methods whose functions often possess specific physical interpretations. Notable examples of statistical point processes encompass the Poisson process Cox (1955), Hawkes process Hawkes (1971), reactive point process Ertekin et al. (2015), self-correcting process Isham and Westcott (1979), and Graphical Event Models Gunawardana and Meek (2016); Bhattacharjya et al. (2018), among others. Conversely, neural point processes Shchur et al. (2021); Chen et al. (2021) exploit the potent learning capabilities of neural networks to model sequences, often outperforming statistical point process methods in prediction performance. These methods are typically predicated on recurrent neural networks and their variants.
2.2 Granger Causality for Event Sequences
Current point process-based Granger causal discovery algorithms for event sequences can be classified into two primary types: the Hawkes process-based methods and the neural point process-based methods. The fundamental assumption of the Hawkes process is that past events stimulate the occurrence of related events in the future. Typical Hawkes process-based methods Zhou et al. (2013); Xu et al. (2016) are focused on designing appropriate kernel functions and regularization techniques. Recently, the THP algorithm Cai et al. (2022) extends the Hawkes process to address the non-i.i.d. issue by employing a topological-temporal convolutional kernel function to model the interdependencies among event sequences. In contrast, neural point process-based methods Xiao et al. (2019); Zhang et al. (2020) do not depend on specific assumptions of generation functions like the Hawkes process. Instead, they employ neural networks to model event sequences, leveraging the expressive power of these networks.
In another vein, some studies model event sequences using time series-based models, such as auto-regressive and constraint-based methods. For instance, Brillinger Brillinger (1994) transforms event sequences into time series, which enables the analysis of event sequences using auto-regressive models. The Granger Causality alignment (GCA) model Li et al. (2021) combines the auto-regressive model with the generation model to learn the Granger causality. PCMCI Runge (2020) extends the PC algorithm Spirtes et al. (2000) to time series data. However, none of these approaches considers the non-i.i.d. event sequences within a topological network.
@@ 2.2 Granger Causality for Event Sequences
3 Granger Causal Discovery using Neural Poisson Auto-Regressive
3.1 Problem Definition
@@
Let an undirected graph $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$ represent the topological network among nodes $\mathbf{N}$, and a directed graph $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$ represent the causal graph among event types $\mathbf{V}$. In this context, $\mathbf{E}_{N}$ signifies the physically connected edges between nodes, and $\mathbf{E}_{V}$ denotes the causal edges between event types. A series of topological event sequences of length $m$, symbolized as $\mathbf{X}=\{(v_{i},n_{i},t_{i})|i\in\{1,\ldots,m\}\}$, are generated by a Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$ within a topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$. Each item in the series consists of $v_{i}\in\mathbf{V}$ representing the event type, $n_{i}\in\mathbf{N}$ representing the topological node, and $t_{i}\in[0,T]$ representing the occurrence timestamp. We can consider the topological event sequences $\mathbf{X}$ as a set of counting processes $\{\mathrm{C}_{v_{i},n_{j}}(t)|v_{i}\in\mathbf{V},n_{j}\in\mathbf{N},t\in[0,T]\}$, where $\mathrm{C}_{v_{i},n_{j}}(t)$ denotes the count of event type $v_{i}$ that has occurred up to $t$ at $n_{j}$. In this work, we divide the continuous interval $[0,T]$ into $T/\Delta$ small intervals, where $\Delta\in\mathbb{R}^{+}$. Within each small interval, we assume that the counting process follows a Poisson process Stoyan et al. (2013).
Based on the above definition, we can extend occurrence numbers of event sequences within the interval $\Delta$, denoted as $\mathbf{O}^{V}=\{O_{t}^{v_{i}}|t\in\{1,\ldots,T/\Delta\},i\in\{1,\ldots,|\mathbf{V}|\}\}$, to occurrence numbers of topological event sequences $\mathbf{O}^{V,N}=\{O_{t}^{v_{i},n_{j}}|t\in\{1,\ldots,T/\Delta\},i\in\{1,\ldots,|\mathbf{V}|\},j\in\{1,\ldots,|\mathbf{N}|\}\}$. Here, $O_{t}^{v_{i},n_{j}}$ represents the number of occurrences of event type $v_{i}$ at node $n_{j}$ within the time interval $(t,t+\Delta]$. The specific value of $\Delta$ can be determined according to the requirements of real-world applications.
As each counting process adheres to a Poisson process in a small interval $\Delta$, the number of occurrences $O^{v_{i},n_{j}}_{t}=\mathrm{C}_{v_{i},n_{j}}(t+\Delta)-\mathrm{C}_{v_{i},n_{j}}(t)$ can be modeled by a Poisson random variable with parameter $\lambda^{v_{i},n_{j}}_{t}$. Thus, the probability of $O^{v_{i},n_{j}}_{t}=o$ is expressed as:
@@
Definition 1 (Causal discovery from topological event sequences).
Given the occurrence numbers of topological event sequences, $\mathbf{O}^{V,N}=\{O_{t}^{v_{i},n_{j}}|t\in\{1,\ldots,T/\Delta\},i\in\{1,\ldots,|\mathbf{V}|\},j\in\{1,\ldots,|\mathbf{N}|\}\}$, and the prior topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$, the goal of causal discovery from topological event sequences is to infer the Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$ among the event types.
In the following, we model the generation process of these topological event sequences by extending an auto-regressive model with the Poisson process to accommodate the non-i.i.d. nature of the data. Furthermore, we devise an amortized inference method to solve the above problem of inferring the latent Granger causal structure.
3.2 Generation Process of the Event Sequences via Topological Neural Poisson Auto-Regressive Model
Figure 2: An illustration of the generation and inference processes for TNPAR. In this figure, $\mathbf{A}_{0:K}$ represents the causal matrices of $\mathcal{G}_{V}$, and $\mathbf{B}_{0:K}$ represents the topological matrices of $\mathcal{G}_{N}$. The solid lines signify the generation process for the data $O_{t}^{v,n}$, while the dashed lines denote the inference process for the causal matrices $\mathbf{A}_{0:K}$.
The generation process of TNPAR is depicted using solid lines in Fig. 2. In this process, the current occurrence number of event type $v_{i}$ at node $n_{j}$ (denoted as $O_{t}^{v_{i},n_{j}}$) is determined by a combination of historical data $\{\mathbf{O}^{V}_{t-1},\mathbf{O}^{V}_{t-2},\dots,\mathbf{O}^{V}_{t-\Omega}\}$, the causal matrices $\mathbf{A}_{0:K}$, and the topological matrices $\mathbf{B}_{0:K}$. Here, $\mathbf{A}_{0:K}$ represents a set of matrices $\{\mathbf{A}_{0},\mathbf{A}_{1},\dots,\mathbf{A}_{K}\}$, where $\mathbf{A}_{k}$ is a $|\mathbf{V}|\times|\mathbf{V}|$ binary matrix that indicates the Granger causality between event types at a geodesic distance Bouttier et al. (2003) of $k$ within $\mathcal{G}_{N}$. Let $A_{k}^{i,j}$ denote the element in row $i$ and column $j$ of $\mathbf{A}_{k}$. Then, if $A_{k}^{i,j}=0$, it implies that, at a geodesic distance of $k$, event type $v_{i}$ does not have a Granger causality with $v_{j}$. Conversely, if $A_{k}^{i,j}=1$, then, at a geodesic distance of $k$, a Granger causality exists from $v_{i}$ to $v_{j}$. Similarly, $\mathbf{B}_{0:K}$ corresponds to another set of matrices $\{\mathbf{B}_{0},\mathbf{B}_{1},\dots,\mathbf{B}_{K}\}$, each of which $\mathbf{B}_{k}$ is a $|\mathbf{N}|\times|\mathbf{N}|$ binary matrix that indicates the physical connections of nodes at a geodesic distance of $k$. Let $B_{k}^{i,j}$ denote the element in row $i$ and column $j$ of $\mathbf{B}_{k}$. If the geodesic distance between node $n_{i}$ and $n_{j}$ is $k$ in $\mathcal{G}_{N}$, then $B_{k}^{i,j}=1$. Otherwise, $B_{k}^{i,j}$ will be set to 00.
@@ 3.2 Generation Process of the Event Sequences via Topological Neural Poisson Auto-Regressive Model
where $\mathbf{O}^{V}_{t-1:t-\Omega}=\{\mathbf{O}^{V}_{t-1},\mathbf{O}^{V}_{t-2},\dots,\mathbf{O}^{V}_{t-\Omega}\}$ represents a set of historical data, $\mathrm{g}(\cdot)$ is a nonlinear function implemented by a neural network that generates data in the $t$-th time interval based on the historical data, and $\epsilon_{t}^{v_{i}}$ denotes the noise term for event type $v_{i}$ in the $t$-th time interval. In order to incorporate the topological relationships of nodes into our model, we extend $\mathbf{O}^{V}_{t-1:t-\Omega}$ to include the topologically connected historical data, using the function $\mathrm{f}^{\mathcal{G}_{N}}_{n_{j}}(\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. This function applies a filter to set the non-topologically connected records for node $n_{j}$ to 00 in the total historical data $\mathbf{O}^{V,N}_{t-1:t-\Omega}$, based on the topological matrices $\mathbf{B}_{0:K}$. This step effectively removes the non-topologically connected information of node $n_{j}$ from $\mathbf{O}^{V,N}_{t-1:t-\Omega}$. Building upon this, our auto-regressive model, incorporating topological information, can be expressed as follows:
@@
Furthermore, to incorporate the Granger causal relationships among event types into the above model, we substitute $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ in Eq. (3) with a function $\mathrm{f}^{\mathcal{G}_{V}}_{v_{i}}(\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{A}_{0:K}\!)$. This function acts as a filter to set the records of event types that do not have a Granger causality with event type $v_{i}$ to 00 in the total historical data $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ based on the causal matrices $\mathbf{A}_{0:K}$. By incorporating both the topological network and the Granger causal structure, we propose the following model for generating topological event sequences:

### v2

@@ Abstract
Learning Granger causality from event sequences is a challenging but essential task across various applications. Most existing methods rely on the assumption that event sequences are independent and identically distributed (i.i.d.). However, this i.i.d. assumption is often violated due to the inherent dependencies among the event sequences. Fortunately, in practice, we find these dependencies can be modeled by a topological network, suggesting a potential solution to the non-i.i.d. problem by introducing the prior topological network into Granger causal discovery. This observation prompts us to tackle two ensuing challenges: 1) how to model the event sequences while incorporating both the prior topological network and the latent Granger causal structure, and 2) how to learn the Granger causal structure. To this end, we devise a unified topological neural Poisson auto-regressive model with two processes. In the generation process, we employ a variant of the neural Poisson process to model the event sequences, considering influences from both the topological network and the Granger causal structure. In the inference process, we formulate an amortized inference algorithm to infer the latent Granger causal structure. We encapsulate these two processes within a unified likelihood function, providing an end-to-end framework for this task. Experiments on simulated and real-world data demonstrate the effectiveness of our approach.
Introduction
@@
Causal discovery from multi-type event sequences is important in many applications. In the realm of intelligent operation and maintenance, identifying the causal structure behind alarm sequences can expedite the location of root causes (Vuković and Thalmann 2022). In social analysis, recovering the causal structure behind users’ behavior sequences can inform the development of effective advertising strategies (Chen et al. 2020).
@@
Lots of methods have been proposed for causal discovery from multi-type event sequences. They fall into two main categories: 1) constraint-based methods, which utilize independence-based tests or measures to estimate the causal structure, and 2) point process-based methods, which employ point processes to model the generation process of event sequences. While the former, including methods such as PCMCI (PC with Momentary Conditional Independence test) (Runge 2020) and transfer entropy-based methods (Chen et al. 2020; Mijatovic et al. 2021), is founded on strict assumptions of the causal mechanism assumptions, the latter, like Hawkes process-based methods (Xu, Farajtabar, and Zha 2016; Zhou, Zha, and Song 2013) and neural point process models (Shchur et al. 2021), is based on the concept of Granger causality (Granger 1969). Given the weaker assumption of Granger causality, identifying the Granger causal structure just by assessing if a sequence is predictive of another, point process-based methods often find preference in real-world applications (Shojaie and Fox 2022) and are therefore the focal point of this work.
@@
Figure 1: An example of the topological event sequences generated by a mobile network in an operation and maintenance scenario. In this context, $\{n_{1},n_{2},n_{3}\}$ represent the network elements, and $\{v_{1},v_{2},v_{3}\}$ denote the event types of the alarms. The term $\mathrm{P}_{v_{1}}(t_{1})$ stands for the distribution of $v_{1}$ alarm at $t_{1}$ timestamp, and so on. Black arrows represent the correct causal edges, whereas the red arrow represents an incorrect edge. (a) Generation process of topological event sequences. The event distribution varies with time and is influenced by both a topological network (connected by the solid lines) and a Granger causal structure (depicted in part (b)). (b) Ground truth of the Granger causal structure. (c) Granger causal structure learned under i.i.d. assumption.
@@
However, these existing point process-based methods are significantly dependent on the i.i.d. assumption to model the generation process of event sequences. In reality, this assumption often falls short due to the complex dependencies that permeate sequences. Consequently, the performance of these methods often leaves much to be desired in real-world applications. For instance, as illustrated in Fig. 1, non-i.i.d. alarm sequences in an operation and maintenance scenario highlight this issue. These event sequences are produced by network elements interconnected by a topological network, indicating the event sequences of different elements are dependent on each other, thereby violating the i.i.d. assumption. In this illustrated scenario, the event type $v_{2}$ in $n_{2}$ is only caused by the event types in topologically connected elements (specifically, $v_{1}$ in $n_{2}$). Existing methods would incorrectly identify the edge $v_{3}\rightarrow v_{2}$ because they erroneously treat event sequences on different nodes as independent. By considering the prior topological network, we can correctly learn that $v_{2}$ is actually caused by $v_{1}$ and discern the true causal structure. Therefore, it is vital to exploit the topological network to model the dependencies among non-i.i.d. event sequences and accurately recover the causal structure from them.
@@
This gives rise to two ensuing challenges: 1) how to explicitly model the generation process of topological event sequences, incorporating both the prior topological network and the latent causal structure, and 2) how to learn the latent Granger causal structure from topological event sequences. To tackle these challenges, Cai et al. (Cai et al. 2022) propose the topological Hawkes process (THP), which extends the time-domain convolution property of the Hawkes process to the time-graph domain. Nevertheless, THP’s applicability in real-world scenarios remains limited due to the complex generation process and distribution of real-world scenarios, which are often too intricate to be accurately represented by the Hawkes process. Hence, it is crucial to devise a distribution-free generation model for non-i.i.d. event sequences that can incorporate both the topological network and the latent Granger causal structure.
@@
In response to these challenges, we develop a unified Topological Neural Poisson Auto-Regressive (TNPAR) model, comprising a generation process and an inference process. In the generation process, we introduce a variant of the Poisson Auto-Regressive model to depict the generation process of topological event sequences. By extending the Auto-Regressive model with Poisson Process and incorporating both the topological network and the Granger causal structure, TNPAR effectively overcomes the non-i.i.d. problem. By representing the distribution using a series of Poisson processes with varying parameters, TNPAR displays enhanced flexibility compared to existing methods that employ a single fixed distribution for the entire sequence. In the inference process, we design an amortized inference algorithm (Zhang et al. 2018; Gershman and Goodman 2014) to effectively learn the Granger causal structure. By integrating these generation and inference processes, we establish a unified likelihood function for TNPAR model, which can be optimized within an end-to-end paradigm.
@@
    By incorporating the prior topological network into the generation model, we offer a distribution-free solution to the non-i.i.d. challenge in causal discovery.
@@
    By treating the causal structure as a latent variable, we devise an amortized inference method to effectively identify the causal structure among the event types.
Related work
Temporal Point Process
@@
Temporal point processes are stochastic processes utilized to model event sequences. They can be bifurcated into two main categories: statistical point processes and neural point processes. On the one hand, statistical point processes emphasize designing appropriate intensity functions, the parameters of which often carry specific physical interpretations. Notable examples of statistical point processes encompass the Poisson process (Cox 1955), Hawkes process (Hawkes 1971), reactive point process (Ertekin, Rudin, and McCormick 2015), and self-correcting process (Isham and Westcott 1979), among others. On the other hand, neural point processes (Shchur et al. 2021) exploit the potent learning capabilities of neural networks to implement the intensity functions, often outperforming statistical point process methods in prediction performance.
Granger Causality for Event Sequences
There are many types of methods for Granger causal discovery from event sequences. For instance, the Hawkes-process-based methods assume that the past events stimulate the occurrence of related events in the future if and only if the former Granger caused the latter. Typical Hawkes process-based methods (Zhou, Zha, and Song 2013; Xu, Farajtabar, and Zha 2016) focus on designing appropriate intensity functions and regularized techniques. Recently, the THP algorithm (Cai et al. 2022) extends the Hawkes process to address the non-i.i.d. issue.Other methods like Graphical Event Models (GEMs) (Bhattacharjya, Subramanian, and Gao 2018) formalize Granger causality using process independence. In contrast, neural point process-based methods (Xiao et al. 2019; Zhang et al. 2020) do not depend on specific assumptions of generation functions like the Hawkes process or GEMs. Instead, they employ neural networks to model event sequences with Granger causality. In another vein, our work is also related to the Granger causal discovery from time series. For instance, Brillinger (Brillinger 1994) aggregates event sequences into time series, which enables the analysis of event sequences using auto-regressive models. The Granger Causality alignment (GCA) model (Li et al. 2023) combines the auto-regressive model with the generation model to learn the Granger causality. Amortized Causal Discovery (ACD) (Löwe et al. 2022) trains an amortized model to infer causal graphs from time series. In addition, we provide a comparison of our work with related methods in Appendix A.
@@
Granger Causal Discovery Using Neural Poisson Auto-Regressive
Problem Definition
@@
Let an undirected graph $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$ represent the topological network among nodes $\mathbf{N}$, and a directed graph $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$ represent the Granger causal graph among event types $\mathbf{V}$. Here, $\mathbf{E}_{N}$ signifies the physically connected edges between nodes, and $\mathbf{E}_{V}$ denotes the causal edges between event types. In this context, an event could trigger its subsequent events in its own node and its topologically connected nodes. A series of topological event sequences of length $m$, symbolized as $\mathbf{X}=\{(v_{i},n_{i},t_{i})|i\in\{1,\ldots,m\}\}$, are generated by a Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$ within a topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$. Each item in the series consists of $v_{i}\in\mathbf{V}$ representing the event type, $n_{i}\in\mathbf{N}$ representing the topological node, and $t_{i}\in[0,T]$ representing the occurrence timestamp.
We can consider the topological event sequences $\mathbf{X}$ as a set of counting processes $\{\mathrm{C}_{v_{i},n_{j}}(t)|v_{i}\in\mathbf{V},n_{j}\in\mathbf{N},t\in[0,T]\}$, where $\mathrm{C}_{v_{i},n_{j}}(t)$ denotes the count of event type $v_{i}$ that has occurred up to $t$ at $n_{j}$. In this work, we divide the continuous interval $[0,T]$ into $\lceil T/\Delta\rceil$ small intervals, where $\Delta\in\mathbb{R}^{+}$ can be determined according to real-world applications. Then, the occurrence numbers of the counting processes can be denoted as $\mathbf{O}^{V,N}=\{O_{t}^{v_{i},n_{j}}|t\in\{1,\ldots,\lceil T/\Delta\rceil\},i\in\{1,\ldots,|\mathbf{V}|\},j\in\{1,\ldots,|\mathbf{N}|\}\}$, where $O_{t}^{v_{i},n_{j}}=\mathrm{C}_{v_{i},n_{j}}(t\times\Delta)-\mathrm{C}_{v_{i},n_{j}}((t-1)\times\Delta)$ represents the occurrence number of $v_{i}$ at $n_{j}$ within $((t-1)\times\Delta,t\times\Delta]$. Within each interval, we assume the counting process follows a Poisson process (Stoyan et al. 2013) with parameter $\lambda^{v_{i},n_{j}}_{t}$. Thus, the probability of $O^{v_{i},n_{j}}_{t}$ can be expressed as follows:
@@
Definition 1 (Granger causal discovery from topological event sequences).
Given the occurrence numbers of topological event sequences, $\mathbf{O}^{V,N}=\{O_{t}^{v_{i},n_{j}}|t\in\{1,\ldots,\lceil T/\Delta\rceil\},i\in\{1,\ldots,|\mathbf{V}|\},j\in\{1,\ldots,|\mathbf{N}|\}\}$, and the prior topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$, the goal of Granger causal discovery from topological event sequences is to infer the Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$ among the event types.
In the following, we propose the TNPAR model to solve the above problem, comprising a generation process and an inference process.
Generation Process of the Event Sequences via Topological Neural Poisson Auto-Regressive Model
Figure 2: An illustration of the generation and inference processes for TNPAR. In this figure, $\mathbf{A}_{0:K}$ represents the causal matrices of $\mathcal{G}_{V}$, and $\mathbf{B}_{0:K}$ represents the topological matrices of $\mathcal{G}_{N}$. Solid lines signify the generation process for the data $O_{t}^{v,n}$, and dashed lines correspond to the inference process for the causal matrices $\mathbf{A}_{0:K}$. Observed variables are denoted by the solid circles, with the latent variables represented by the dashed circles.
TNPAR’s generation process is depicted using solid lines in Fig. 2. In this process, $O_{t}^{v_{i},n_{j}}$ is determined by a combination of historical data $\{\mathbf{O}^{V,N}_{t-1},\mathbf{O}^{V,N}_{t-2},\dots,\mathbf{O}^{V,N}_{t-\Omega}\}$, the causal matrices $\mathbf{A}_{0:K}$, and the topological matrices $\mathbf{B}_{0:K}$. Here, $\mathbf{A}_{0:K}$ represents a set of matrices $\{\mathbf{A}_{0},\mathbf{A}_{1},\dots,\mathbf{A}_{K}\}$, where $\mathbf{A}_{k}$ is a $|\mathbf{V}|\times|\mathbf{V}|$ binary matrix that indicates the Granger causality between event types at a geodesic distance (Bouttier, Di Francesco, and Guitter 2003) of $k$ within $\mathcal{G}_{N}$. Let $A_{k}^{i,j}$ denote the element in row $i$ and column $j$ of $\mathbf{A}_{k}$. Then, if $A_{k}^{i,j}=0$, it implies that, at a geodesic distance of $k$, event type $v_{i}$ does not have a Granger causality with $v_{j}$. Otherwise, $A_{k}^{i,j}$ will be $1$. Similarly, $\mathbf{B}_{0:K}$ corresponds to another set of matrices $\{\mathbf{B}_{0},\mathbf{B}_{1},\dots,\mathbf{B}_{K}\}$, each of which $\mathbf{B}_{k}$ is a $|\mathbf{N}|\times|\mathbf{N}|$ binary matrix that indicates the physical connections of nodes at a geodesic distance of $k$. Let $B_{k}^{i,j}$ denote the element in row $i$ and column $j$ of $\mathbf{B}_{k}$. If the geodesic distance between node $n_{i}$ and $n_{j}$ is $k$ in $\mathcal{G}_{N}$, then $B_{k}^{i,j}=1$. Otherwise, $B_{k}^{i,j}$ will be set to 00.
@@
where $\mathbf{O}^{V}_{t-1:t-\Omega}=\{\mathbf{O}^{V}_{t-1},\mathbf{O}^{V}_{t-2},\dots,\mathbf{O}^{V}_{t-\Omega}\}$ represents a set of historical data, $\mathrm{g}(\cdot)$ is a nonlinear function implemented by a neural network that generates data in the $t$-th time interval based on the historical data, and $\epsilon_{t}^{v_{i}}$ denotes the noise term for event type $v_{i}$ in the $t$-th time interval. In order to incorporate the topological relationships of nodes into our model, we extend $\mathbf{O}^{V}_{t-1:t-\Omega}$ to include the topologically connected historical data, using the function $\mathrm{f}^{\mathcal{G}_{N}}_{n_{j}}(\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{B}_{0:K})$. This function applies a filter to set the non-topologically connected data for node $n_{j}$ to 00 in $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ based on $\mathbf{B}_{0:K}$, which effectively removes the non-topologically connected information in the generation process of $O_{t}^{v_{i}}$. Building upon this, our auto-regressive model, incorporating topological information, can be expressed as follows:
@@
Furthermore, to incorporate the Granger causal structure into the above model, we substitute $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ in Eq. (3) with a function $\mathrm{f}^{\mathcal{G}_{V}}_{v_{i}}(\mathbf{O}^{V,N}_{t-1:t-\Omega},\mathbf{A}_{0:K}\!)$. This function acts as a filter to set the historical data that do not have a Granger causality with $v_{i}$ to 00 in $\mathbf{O}^{V,N}_{t-1:t-\Omega}$ based on $\mathbf{A}_{0:K}$. By incorporating both the topological network and the Granger causal structure, we propose the following model for generating topological event sequences: